OpenStudy (notamathgenius):

A trapeze artist swings in simple harmonic motion with a period of 3.8s. Calculate the length of the cables supporting the trapeze.

OpenStudy (irishboy123):

pendulum \[T = 2 \pi \sqrt{\frac{l}{g}}\]